Consider a spherical blackbody of temperature T=230 oC and radius R=0.20m that radiates energy to an environment of very low temperature. Suppose you surround the sphere with a black, thin, spherical shell of radius 6R. What is the temperature of the shell in Celsius when it reaches a steady state? Remember the inner surface of the large sphere is always absorbing energy from the black body and the outer surface is loosing energy to the very low temperature surroundings.
  oC
Answer #1

Pin = Pout Р= ТАТУ lin= t yor?:( 230 + 273-1574 Pout = 0 4 0 COR?.T4 at steady estate Pin = Pout (230 +273-1539 = 6? T4 T 840